OpenStudy (aravindg):

solve sec theta-cosec theta=4/3

OpenStudy (anonymous):

\[sec \theta -cosec \theta = \frac{4}{3}\]

OpenStudy (aravindg):

.............

OpenStudy (anonymous):

Conversion into sin and cos will be good

OpenStudy (aravindg):

show

OpenStudy (aravindg):

............

OpenStudy (anonymous):

\[\frac{sin\theta - cos\theta}{sin\theta *cos\theta}=\frac{4}{3}\]

OpenStudy (anonymous):

Are you sure this is the question ?

OpenStudy (aravindg):

yes

OpenStudy (aravindg):

vry bad

OpenStudy (anonymous):

could you check the given data again?